OpenStudy (anonymous):

Solve the system using elimination? -2x+2y+3z=0, -2x-y+z=-3, 2x+3y+3z=5

OpenStudy (campbell_st):

well looking at the equations, which letter do you think you should eliminate 1st..?

OpenStudy (anonymous):

x?

OpenStudy (anonymous):

Answer is x = 1, y = 1 and z = 0.

OpenStudy (campbell_st):

so add equations 1 and 3, this will eliminate x -2x + 2y + 3x = 0 2x + 3y + 3z = 5 -----------------\ 5y + 6z =5 equation 4 same for equation 2 and 3 -2x - y + z = -3 2x + 3y + 3z = 5 ------------------ 2y + 4z = 2 divide each term by 2 and you get y + 2z = 1 equation 5 so you now need to solve 5y + 6z = 5 y + 2z = 1 so multiply equation 5 by -3 and add to equation 4 5y + 6z = 5 -3y - 6z = -3 ---------------- 2y = 2 solve for y. Find z by substituting into equations 4 or 5 then when you gte that you can find x by substituting your y and z values into any of the initial equations. hope it helps